OpenStudy (anonymous):

Combustion of a 10.16 mg sample of methanol produces 11.43 mg of water and 13.96 mg of carbon dioxide. Molecular weight measurements show the molecular weight of methanol to be 32 g/mole. Show that these analytical data lead to the molecular formula CH4O for methanol.

OpenStudy (anonymous):

It sounds like what you need to do is show that the given data adhere to the ratios given by the combustion reaction, \[4CH_4O+6O_2~~\implies~~8H_2O+4CO_2\]

OpenStudy (anonymous):

The equation above gives you the ratios in which moles of a given substance occur in the reaction, i.e. 4 moles of methanol will yield 4 moles of carbon dioxide. Molar masses of reactants/products: \[\begin{array}{c|c} \text{element}&\text{molecular mass }\bigg(\text{approximate, }\dfrac{\text{g}}{\text{mol}}\bigg)\\ \hline CH_4O&12+4(1)+16=32\\ O_2&2(16)=32\\ H_2O&2(1)+16=18\\ CO_2&12+2(16)=44 \end{array}\] Using this info, you should be able to compute how many moles of reactants and products you obtained from the reaction. Finally, check the ratios. If they're similar to the ones in the equation, then the data aligns with the model.